# PixWizard: Versatile Image-to-Image Visual Assistant with Open-Language Instructions

### v2

@@ 4.4 Ablation Study
In this section, we performed ablation studies to assess the impact of each component’s design and the training process on learning in PixWizard. Given computational limitations, we conducted the ablation on PixWizard, training it for 40k steps.
@@
Structural-Aware vs. Semantic-Aware Guidance. We demonstrate the importance of incorporating both structure-aware and semantic-aware guidance to boost PixWizard’s performance across a variety of heterogeneous tasks. To illustrate this, we trained two additional models: $(M1)$ using only the structure-aware module, and $(M2)$ using only the semantic-aware module. As shown in Table 4, the $M1$ model performs better on tasks that require preserving overall image structure and details. In contrast, the $M2$ model, which relies solely on cross-attention for injecting image features, struggles with most visual tasks, often generating images that deviate significantly from the input. However, as noted by other methods (Ye et al., 2023; Hu et al., 2024), semantic guidance typically excels in tasks that require more flexibility, such as text-to-image generation, conditional generation, and image editing. Compared to PixWizard, which integrates both modules, the combined approach enables the model to handle a wider variety of tasks, achieving balanced performance across different scenarios.
Influence of Dynamic Semantic Tokens Sampling. As shown in Table 4, the impact of using Dynamic Semantic Tokens Sampling (DSTS) on task performance is minimal. However, on average, using the DSTS method leads to better overall performance. This suggests that it’s unnecessary to model semantic tokens for the entire image. Since different tasks focus on different semantic regions, dynamically sampling the relevant features allows the model to handle each task more efficiently. Additionally, using fewer tokens reduces the computational load during attention calculations, leading to faster inference times.
@@ 5 Discussion and Conclusion
In this work, we explore how to build a versatile interactive image-to-image visual assistant from three key aspects: task definition, data construction, and model architecture. Our goal is to create a system that can precisely follow free-form user instructions for image generation, manipulation, and translation. Our approach, PixWizard, eliminates the need for task-specific design choices and achieves highly competitive performance across a diverse set of tasks, with strong generalization capabilities.
However, this work has some limitations. First, the current model architecture does not yet support multi-image input conditions, which is an increasingly important and valuable research area. Second, there is still room for improvement, particularly in challenging tasks such as segmentation and image grounding when compared to specialized models. Additionally, the performance of the text encoder and foundation model plays a crucial role. Better text encoding methods enable the model to more accurately understand and execute human instructions, while larger and more robust model architectures directly improve the quality of the final output. Notably, the modules and strategies we propose in PixWizard can be easily applied to other powerful text-to-image generators. In the future, we will explore using more advanced diffusion foundation models, such as SD3 and FLUX, and continue to push this promising direction forward until we achieve the vision field’s "GPT-4 moment."
@@ Appendix A Related Work
Diffusion Models. Diffusion models estimate the data distribution by modeling the gradient of the noise-perturbed data distributions (Sohl-Dickstein et al., 2015; Song & Ermon, 2019; Ho et al., 2020; Dhariwal & Nichol, 2021; Song et al., 2021). They have demonstrated remarkable performance in various fields, ranging from text-to-image generation (Rombach et al., 2022a; Saharia et al., 2022; Betker et al., 2023; Jiang et al., 2024a; Zhang et al., 2023b), controllable generation (Zhang et al., 2023a; Ye et al., 2023), and image editing (Avrahami et al., 2022; Brooks et al., 2023; Kawar et al., 2023) to video (Ho et al., 2022; Brooks et al., 2024), audio (Kong et al., 2021; Huang et al., 2023), 3D (Guo et al., 2023; 2024), and motion (Tevet et al., 2023; Zhang et al., 2024b) generation. Beyond generation, recent works have also exhibited diffusion models’ capabilities in computer vision tasks, such as semantic segmentation (Baranchuk et al., 2022; Xu et al., 2023), depth estimation (Ke et al., 2024b; Lee et al., 2024), and image restoration (Xia et al., 2023). Benefiting from the visual knowledge learned from large-scale pretraining, these works open up the potential for adapting pretrained diffusion models to downstream tasks in a generative manner, with excellent capabilities like handling inherent uncertainties and zero-shot generalization.
@@ B.2 Image Grounding
User Sketch to Image. Following ControlNet (Zhang et al., 2023a), we generate human sketches from images by applying HED boundary detection (Xie & Tu, 2015) combined with strong data augmentations, including random thresholds, random masking of sketch portions, random morphological transformations, and random non-maximum suppression. This process results in 0.5 million sketch-image-caption pairs.
@@
Normal Map to Image. The surface normals are generated using DSINE Bae & Davison (2024), which contributed to the depiction of surface orientation and texture details in the generated images. Finally, we obtain 0.8M normal-image-caption pairs.
@@ B.4 Dense Image Prediction
Handling free-form user prompts, rather than fixed task-specific prompts, increases PixWizard’s usability as a general visual assistant. In Fig. 13, we present examples of task templates rephrased by GPT-4o.
@@
Appendix C More Detailed Descriptions for the Architecture
@@ C.1 Preliminaries
C.2 Classifier-free Guidance for Multi-Modal Conditions
@@
C.3 Inference Efficiency with Fewer Steps
Compared to diffusion-based methods (Gan et al., 2024; Hu et al., 2024; Wang et al., 2023c), our model demonstrates better inference efficiency under a low number of function evaluations (NFE) regime. By leveraging the linear interpolation schedule with flow matching objective, PixWizard can learn straighter trajectories, thus requiring fewer sampling steps to achieve decent performance. Drawing inspiration from the observation that flow-based models exhibit higher discretization error at the beginning of generation, we employ the shifted time schedule which allocates more steps near noise. By further integrating high-order ODE solvers with improved schedules, such as Heun and Midpoint, our model produces reasonable output using only 30-60 NFEs, compared to 100 NFEs used in diffusion-based counterparts.
@@ Appendix D More Experimental Results
D.1 Image Restoration
@@ D.2 Image Grounding
PixWizard 34.17 46.44 41.25 43.43 36.49 35.30 38.93
Results. Table 6 reports the results for referring segmentation. Our model demonstrates strong performance under two different mask extraction methods, outperforming InstructDiffusion across almost all evaluation datasets. In the various test sets of RefCOCO and RefCOCO+, our performance is comparable to Unified-IO, but on gRefCOCO, we significantly outperform Unified-IO (34.17 vs. 17.31). However, it is important to note that these unified methods still lag considerably behind those specifically designed for referring segmentation. Additionally, we observed that using a generative model for grounding tasks presents a significant challenge due to the model’s limited perception and localization capabilities. This often results in failures to correctly locate objects when multiple targets are present, highlighting the need for improvements in instruction and understanding in future work.

### v3

@@ 4.4 Ablation Study
We performed ablation studies to assess the impact of each component’s design and the training process on learning in PixWizard. Given computational limitations, we conducted the ablation on PixWizard, training it for 40k steps.
@@
Structural-Aware vs. Semantic-Aware Guidance. We demonstrate the importance of integrating both structure-aware and semantic-aware guidance to enhance PixWizard’s performance across diverse tasks. To validate this, we trained two additional models: $M1$ with only the structure-aware module and $M2$ with only the semantic-aware module. Results show that $M1$ performs better on tasks requiring preservation of image structure and details, while $M2$, which relies solely on cross-attention for injecting image features, struggles with most visual tasks and often generates outputs that deviate from the input. However, as noted by other methods (Ye et al., 2023; Hu et al., 2024), semantic guidance excels in tasks needing flexibility, such as text-to-image generation, conditional generation, and image editing. In comparison, PixWizard, which combines both modules, achieves balanced performance across a broader range of tasks.
Influence of Dynamic Semantic Tokens Sampling. As shown in Table 4, the impact of Dynamic Semantic Tokens Sampling (DSTS) on task performance is minimal, but it improves overall performance on average. This indicates that modeling semantic tokens for the entire image is unnecessary. By dynamically sampling relevant features for each task’s focus, the model operates more efficiently. Moreover, using fewer tokens reduces computational load during attention calculations, resulting in faster inference.
@@ 5 Discussion and Conclusion
In this work, we explore how to build a versatile interactive image-to-image visual assistant from three key aspects: task definition, data construction, and model architecture. Our goal is to create a system that can precisely follow free-form user instructions for image generation, manipulation, and translation. Our PixWizard eliminates the need for task-specific design choices and achieves highly competitive performance across a diverse set of tasks, with strong generalization capabilities.
However, this work has some limitations. First, the current model architecture does not yet support multi-image input conditions, which is an increasingly important research area. Second, there is room for improvement in challenging tasks like segmentation and image grounding compared to specialized models. Additionally, the performance of the text encoder and foundation model also plays a crucial role. Better text encoding improves the model’s ability to understand and execute instructions, while larger, more robust architectures enhance output quality. Notably, the modules and strategies proposed in PixWizard can be easily applied to other powerful text-to-image generators. In the future, we plan to explore more advanced diffusion models, such as SD3 and FLUX, and continue advancing this field toward a "GPT-4 moment".
6 Acknowledgements
We thank Ziheng Wu and Tonton Su for their valuable comments and suggestion to this project, including data construction, model discussions, and evaluation work. This project is funded in part by National Key R&D Program of China Project 2022ZD0161100, by the Centre for Perceptual and Interactive Intelligence (CPII) Ltd under the Innovation and Technology Commission (ITC)’s InnoHK, by NSFC-RGC Project N_CUHK498/24. Hongsheng Li is a PI of CPII under the InnoHK.
Appendix
@@ Appendix A Related Work
Diffusion Models. Diffusion models estimate the data distribution by modeling the gradient of the noise-perturbed data distributions (Sohl-Dickstein et al., 2015; Song & Ermon, 2019; Ho et al., 2020; Dhariwal & Nichol, 2021; Song et al., 2021). They have demonstrated remarkable performance in various fields, ranging from text-to-image generation (Rombach et al., 2022a; Saharia et al., 2022b; Betker et al., 2023; Jiang et al., 2024a; Zhang et al., 2023b), controllable generation (Zhang et al., 2023a; Ye et al., 2023), and image editing (Avrahami et al., 2022; Brooks et al., 2023; Kawar et al., 2023) to video (Ho et al., 2022; Brooks et al., 2024), audio (Kong et al., 2021; Huang et al., 2023), 3D (Guo et al., 2023; 2024), motion (Tevet et al., 2023; Zhang et al., 2024b) generation, and reasoning enhancement (Guo et al., 2025). Beyond generation, recent works have also exhibited diffusion models’ capabilities in computer vision tasks, such as semantic segmentation (Baranchuk et al., 2022; Xu et al., 2023), depth estimation (Ke et al., 2024b; Lee et al., 2024), and image restoration (Xia et al., 2023). Benefiting from the visual knowledge learned from large-scale pretraining, these works open up the potential for adapting pretrained diffusion models to downstream tasks in a generative manner, with excellent capabilities like handling inherent uncertainties and zero-shot generalization.
@@ B.2 Image Grounding
User Sketch to Image. Following ControlNet (Zhang et al., 2023a), we generate human sketches from images by applying HED boundary detection (Xie & Tu, 2015) combined with strong data augmentations, including random thresholds, random masking of sketch portions, random morphological transformations, and random non-maximum suppression. This process results in 0.4 million sketch-image-caption pairs.
@@
Normal Map to Image. The surface normals are generated using DSINE (Bae & Davison, 2024), which contributed to the depiction of surface orientation and texture details in the generated images. Finally, we obtain 0.8M normal-image-caption pairs.
@@ B.4 Dense Image Prediction
Handling free-form user prompts, rather than relying on fixed task-specific prompts, significantly enhances PixWizard’s flexibility and usability as a general visual assistant. In Fig. 10, we showcase various examples where task-specific templates are dynamically rephrased and adapted by GPT-4o
@@
Appendix C More Details for the Architecture
@@ C.1 Preliminaries
C.2 Task Embedder
As shown in Fig. 11, introducing the task embedder helps adaptively cluster similar task instructions in the latent space while separating those from different tasks, guiding the model’s generation process in the correct direction.
Figure 11: t-SNE visualization of the global text embeddings. Each dot represents a user instruction.
C.3 Classifier-free Guidance for Multi-Modal Conditions
@@
In Fig. 12, we illustrate the effects of the two parameters on the generated samples. As shown, changes in both $W_{I}$ and $W_{T}$ significantly influence the results. Specifically, in PixWizard, $W_{I}$ strongly affects the color distribution of the generated images. An increase in $W_{I}$ leads to more saturated and unrealistic colors. This can also negatively impact tasks such as depth estimation, where subtle depth variations may be represented as solid black regions. In contrast, the effects of $W_{T}$ vary depending on the task. For example, in dense image prediction tasks, increasing $W_{T}$ can degrade the quality of the estimation. However, for generative tasks, a higher $W_{T}$ typically has little effect.
Figure 12: Impact of classifier-free guidance.
C.4 Any Resolution
PixWizard inherits the dynamic partitioning and padding scheme introduced by (Zhuo et al., 2024), allowing the model to handle images of any resolution and aspect ratio during both fine-tuning and inference.
The dynamic partitioning and padding scheme is introduced to optimize image token handling, avoiding fixed resolutions. Specifically, given constraints on the maximum number of patches and the maximum aspect ratio, the method defines a set of candidate patch partitions. For each input image size, it determines the optimal patch partition based on a matching ratio between the input size and the candidate partition, selecting the best fit. The input image is then resized accordingly using the chosen partition. This dynamic process leads to varying sequence lengths of patch tokens, so padding is applied to align the lengths across batches using a pad token. To prevent unwanted interactions between pad and regular tokens, attention masks are also introduced. This approach maximizes efficiency while preserving image integrity during embedding.
However, in practice, the required resolutions for different tasks can vary significantly. To support more flexible handling of arbitrary resolutions while preserving the original resolution as much as possible, we use $[512^{2},768^{2},1024^{2}]$ as resolution centers to generate candidate patch partitions. During training, we group data with similar resolutions into buckets, ensuring that sequence lengths within each batch are comparable, minimizing padding tokens, and improving training efficiency. During inference, by incorporating NTK-Aware Scaled RoPE (Peng & Quesnelle, 2023) and sandwich normalization, PixWizard demonstrates exceptional resolution extrapolation.
@@ Appendix D More Experimental Results
D.1 Implementation Details
Dense Image Prediction. For semantic segmentation, we assign labels by identifying the nearest neighbor RGB color value, and accuracy is evaluated using the Mean Intersection over Union (mIoU) metric. For monocular depth estimation, we average the output image across the three channels and apply the inverse of the linear transformation used during training to obtain depth estimates within the range of $[0,10]$ meters. Accuracy is evaluated using the Root Mean Square Error (RMSE). For surface normal estimation, we recover the corresponding normal vectors from the output image and use the Mean Angle Error to assess accuracy.
Controllable Generation. We mainly evaluated PixWizard’s ability based on two conditions: canny edge maps and depth maps. Following ControlNet++ (Li et al., 2024), we measured controllability using RMSE for depth maps and F1-Score for canny edges, comparing input conditions with features from the generated images. Image quality and text alignment were assessed using FID (Fréchet Inception Distance) and CLIP-Score. All experiments were conducted at a resolution of $512\times 512$.
Image Inpainting. We used latent diffusion (Rombach et al., 2022b) to measure FID and LPIPS, focusing on samples where 40-50% of the image area was inpainted. For outpainting, we followed MaskGIT (Chang et al., 2022a) settings, extending the image by 50% and evaluating performance with FID and Inception Score (IS) on $512\times 512$ crops from the Places dataset (2017a).
Text-to-image Generation. We evaluated PixWizard with two methods: visual examples and automatic metrics, including the Human Preference Score (HPS) v2 (Wu et al., 2023) and Zero-shot FID-30K on the MS-COCO dataset (Lin et al., 2014).
D.2 Image Restoration
@@ D.3 Image Grounding
PixWizard 29.07 44.38 40.13 42.09 36.49 35.30 38.93
Results. Table 6 reports the results for referring segmentation. Our model demonstrates strong performance under two different mask extraction methods, outperforming InstructDiffusion across almost all evaluation datasets. In the various test sets of RefCOCO and RefCOCO+, our performance is comparable to Unified-IO, but on gRefCOCO, we significantly outperform Unified-IO (33.65 vs. 17.31). However, it is important to note that these unified methods still lag considerably behind those specifically designed for referring segmentation. Additionally, we observed that using a generative model for grounding tasks presents a significant challenge due to the model’s limited perception and localization capabilities. This often results in failures to correctly locate objects when multiple targets are present, highlighting the need for improvements in instruction and understanding in future work.
D.4 More Precise Editing
Based on the robust inpainting and grounding capability of PixWizard, we find that it allows user for more precise image editing tasks: (i) Remove Anything. It tackles the object removal problem (Criminisi et al., 2003; 2004; Elharrouss et al., 2020), enabling users to seamlessly remove specific objects. As shown in Fig. 13, PixWizard first generates a target mask based on user instructions, then fills the area with appropriate background details. (ii) Replace Anything. It lets users swap objects in an image, replacing them with specified objects while maintaining background consistency. (iii) Add Anything. This allows users to insert objects into an image. By adding a mask and providing a text prompt, PixWizard generates the desired content using its advanced inpainting ability.
Figure 13: Visualization results of Remove, Replace and Add Anything.
D.5 Multi-Hot Gumbel-Softmax
In the implementation of the Multi-Hot Gumbel-Softmax (MHGS), the pseudocode is defined as follows:
def MHGS(logits, temp=1, dim=-1, sample_tokens=16):
# Add Gumbel noise and scale by temperature
gumbels = (logits + GumbelNoise(shape=logits.shape)) / temp
# Apply Softmax to obtain soft outputs
y_soft = Softmax(gumbels, dimension=dim)
# Select top-k values for discrete output
indices = Top-K(y_soft, k=sample_tokens, dimension=dim)
# Create a hard multi-hot tensor from indices
y_hard = MultiHotTensor(indices, shape=logits.shape)
# Combine hard and soft outputs while preserving gradients
ret = y_hard - StopGradient(y_soft) + y_soft
return ret